# biblio.bib

### v2

@@ 1.1 Locally optimal cuts
We say that a cut is locally optimal if for every node at least half of its incident edges are cut edges. Put otherwise, changing the label of $v$ from $\ell(v)$ to $-\ell(v)$ does not increase the size of the cut. Such a cut always exists and is trivial to find in the centralized setting: start with any labeling, find any node that violates the constraint (we call these unhappy nodes) and change its label, and repeat until all nodes are happy. Note that changing the label of an unhappy node $v$ will make node $v$ happy but it might make some of its neighbors unhappy. However, this process has to converge after at most |$\absolutevalue{E}$ | steps, since the size of the cut increases in each step by at least one.
@@
While it is trivial to find a locally optimal cut with a centralized algorithm, in this work we are interested in distributed algorithms for finding a locally optimal cut. We work in the usual LOCAL model [linial-1992-locality-in-distributed-graph-algorithms, peleg-2000-distributed-computing-a-locality-sensitive] of distributed computing: Each node $v\in V$ represents a computer and each edge $\{u,v\}$ represents a communication link between two computers. Initially each node is only aware of its own unique identifier, its degree, and the total number of nodes $n$. Computation proceeds in synchronous rounds; in each round each node can send a message to each neighbor, receive a message from each neighbor, and update its own state. Eventually each node $v$ has to stop and output its own part of the solution, here $\ell(v)$. The running time of the algorithm is $T(n)$ if in any $n$-node graph all nodes stop after at most $T(n)$ rounds.
@@
[…]
Most natural LCL problems are by now very well-understood, especially for the deterministic LOCAL model. Indeed, for some restricted families of LCLs and restricted families of graphs, there is now a complete classification of the round complexity of all such problems, see e.g. [balliu-brandt-etal-2020-classification-of-distributed, chang-studeny-suomela-2023-distributed-graph-problems, brandt-hirvonen-etal-2017-lcl-problems-on-grids, balliu-brandt-etal-2022-efficient-classification-of, balliu-brandt-etal-2019-the-distributed-complexity-of]. For example, there is a broad class of symmetry-breaking problems, all of which have round complexity $\Theta(\operatorname{log}^{*}n)$ [cole-vishkin-1986-deterministic-coin-tossing-with, goldberg-plotkin-shannon-1988-parallel-symmetry, linial-1992-locality-in-distributed-graph-algorithms, naor-1991-a-lower-bound-on-probabilistic-algorithms-for], and many other problems [ghaffari-hirvonen-etal-2020-improved-distributed-degree] have turned out to be asymptotically as hard as the sinkless orientation problem [brandt-fischer-etal-2016-a-lower-bound-for-the], which has round complexity $\Theta(\operatorname{log}n)$ in the deterministic LOCAL model [brandt-fischer-etal-2016-a-lower-bound-for-the, chang-kopelowitz-pettie-2019-an-exponential-separation, ghaffari-su-2017-distributed-degree-splitting-edge].
Locally optimal cut is perhaps the most elementary LCL problem whose distributed complexity is currently wide open. We do not even know if it admits any nontrivial algorithm with round complexity $o(n)$, not even for the smallest nontrivial degree bound $\Delta=3$. (The case of $\Delta=2$ is fully understood, as it is in essence equivalent to the weak-$2$-coloring problem, which can be solved in $O(\operatorname{log}^{*}n)$ rounds, and this is also tight, see e.g. [balliu-hirvonen-etal-2019-hardness-of-minimal-symmetry].)
@@
What makes locally optimal cuts particularly intriguing is that many commonly-used lower bound techniques fail. In particular, one cannot use existential graph-theoretic arguments (in the spirit of e.g. the lower bound for coloring trees in [linial-1992-locality-in-distributed-graph-algorithms] or the lower bound for approximate coloring in [coiteux-roy-d-amore-etal-2024-no-distributed-quantum]), since a locally optimal cut always exists in any graph.
It is, however, possible to construct a nontrivial reduction that connects locally optimal cuts with the sinkless orientation problem [balliu-hirvonen-etal-2019-locality-of-not-so-weak], and this gives the following lower bounds: we need at least $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model, and at least $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model.
This is essentially everything that is known about locally optimal cuts before this work. There are no upper bounds better than the trivial $O(n)$. Furthermore, the lower bounds do not hold in stronger models such as quantum-LOCAL [gavoille-kosowski-markiewicz-2009-what-can-be-observed, arfaoui-fraigniaud-2014-what-can-be-computed-without, akbari-coiteux-roy-etal-2025-online-locality-meets, balliu-brandt-etal-2025-distributed-quantum-advantage]—there we do not even know if the true complexity of the problem is closer to $O(1)$ or $\Omega(n)$.
We note that locally optimal cuts have been studied e.g. in the context of descriptive combinatorics also under the names of unfriendly colorings and unfriendly partitions [shelah-milner-2012-graphs-with-no-unfriendly-partitions, conley-2014-measure-theoretic-unfriendly-colorings, conley-marks-unger-2020-measurable-realizations-of, conley-tamuz-2020-unfriendly-colorings-of-graphs-with].
@@ 1.2 Local potential problems
For example, in the locally optimal cut problem, our node labels are $\{-1,+1\}$. A node is “happy” if at least half of its incident edges are cut edges (this is a local definition in the sense that it only depends on the constant-radius neighborhood of the node). We assign local potential $1$ to each cut edge and potential 00 to all other edges (this is a local potential in the sense that it only depends on the constant-radius neighborhood of the edge). Now we can always turn unhappy nodes into happy nodes by flipping their label, and this will improve the total potential around the node.
@@
and the task is to maximize the sum of local utilities. For example, the maximum cut problem is such an optimization problem; the local utility function assigns value $1$ to each cut edge and 00 to all other edges.
@@
Conversely, super-polylogarithmic lower bounds on the distributed round complexity now directly imply that a problem cannot be represented with the help of a local potential. To give a simple example, stable matchings require a linear number of rounds, even in bounded-degree bipartite graphs [floreen-kaski-etal-2010-almost-stable-matchings-by], so this problem cannot be represented with local potential functions. Results of this form may be of interest also in a much broader context outside distributed computing.
@@
Key examples of local potential problems include defective coloring problems, see e.g. [barenboim-elkin-2013-distributed-graph-coloring, Chapter 6], [chung-pettie-su-2017-distributed-algorithms-for-the, fischer-ghaffari-2017-sublogarithmic-distributed]. Assume that we have $c$ colors and we want to have defect at most $d$, that is, each node can have at most $d$ neighbors with the same color. If $c(d+1)>\Delta$, this is a local potential problem. Here the potential function is the number of properly colored edges. If a node is unhappy (it has more than $d$ neighbors with the same color), it can find a color that makes it happy. Our algorithm can be used to solve this problem in for any constants $c$, $d$ and $\Delta$ with $c(d+1)>\Delta$ in polylogarithmic number of rounds.
@@ 1.3 New techniques and key ideas
Our algorithm uses the MPX clustering algorithm [miller-peng-xu-2013-parallel-graph-decompositions-using, elkin-neiman-2022-distributed-strong-diameter-network] to divide the nodes into polylogarithmic-radius clusters with small boundaries. Each cluster can make nodes far from boundaries safe; we can simply compute for each cluster the result of repeatedly finding for each internal node an improving set $A$, and flipping $A$ until all internal nodes are safe.
@@
Our algorithm is randomized, but it can be derandomized (with polylogarithmic cost) by using the standard techniques [ghaffari-harris-kuhn-2018-on-derandomizing-local, rozhon-ghaffari-2020-polylogarithmic-time-deterministic].
@@ 2.1 Notation and definitions
In this paper, we consider simple undirected graphs unless otherwise stated. For a graph $G=(V,E)$, we denote its set of nodes by $V$ and its set of edges by $E$. If $G$ is not clear from the context we use $V(G)$ to denote the set of nodes of $G$ and $E(G)$ to denote the set of edges of $G$. For any two nodes $u,v\in V(G)$ the distance between $u$ and $v$ is the number of hops in a shortest path between $u$ and $v$, and is denoted by $\operatorname{dist}(u,v)$. For any node $v\in V(G)$, the radius-$r$ neighborhood $\mathcal{N}_{r}[v]$ of $v$ is the set of all nodes at distance at most $r$ from $v$. For any two graphs $G,H$, the intersection between $G$ and $H$ is the graph $G\cap H=(V(G)\cap V(H),E(G)\cap E(H))$, and the union between $G$ and $H$ is the graph $G\cup H=(V(G)\cup V(H),E(G)\cup E(H))$. The difference between two graphs $G$ and $H$ is the graph $G\setminus H=(V(G),E(G)\setminus E(H))$.
@@
In this paper, we consider the LOCAL model of distributed computation [linial-1992-locality-in-distributed-graph-algorithms]. We are given a distributed network of $n$ nodes/processors, each of which is capable of arbitrary (but finite) computation. Nodes are connected by edges which represent communication links, and are able to send messages of arbitrarily large (but finite) size to each other: note that nodes can discriminate between different communication links through ports, which are numbered from $1$ to $\Delta$, where $\Delta$ is the maximum degree of the network, that are adversarially assigned. Each node is also given a unique identifier in the set $[n^{c}]$ for some constant $c\geq 1$, which is, again, adversarially assigned. The communication network is represented by a simple undirected graph $G=(V,E)$ where $V$ is the set of nodes/processors and $E$ is the set of edges/communication links. All nodes run the same algorithm and computation proceeds in synchronous rounds. In each round, each node sends messages to its neighbors, receives messages from all of its neighbors, and performs local computation. The input of a problem is given in the form of a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of labels. The output of a node is an assignment of labels to the node itself and/or to its incident half-edges: note that this information can be stored in the node through the use of port numbers. When a node determines its own output, we say that the node terminates its computation. The running time of a LOCAL algorithm is the number of communication rounds it takes for all nodes to produce their output. We also refer to the running time of an algorithm as the locality of the algorithm. The complexity of a problem is the minimum running time (across all possible algorithms) needed to solve the problem. We can also assume that nodes are equipped with some source of randomness (independently of other nodes), such as an arbitrarily long random bit string, which they can use during the computation. In this case, we talk about the randomized LOCAL model as opposed to the deterministic LOCAL model. In case of randomized computation, we require that the problem is solved with probability at least $1-1/n$.
@@ 2.2 Problems of interest
Let $H=(V,E)$ be any graph, and let $v\in V$ be any node of $H$. The pair $(H,v)$ is said to be a centered graph and $v$ is said to be the center of $(H,v)$. We say that the pair $(H,v)$ has radius $r$ if all nodes in $H$ are at distance at most $r$ from $v$ and all edges in $H$ have at least one endpoint at distance at most $r-1$ from $v$.
@@
In the following, we define what it means to satisfy a set of constraints. For a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ and a node $v\in V$, we denote by $G_{r}(v)$ the $(\mathcal{V},\mathcal{E})$-labeled subgraph of $G$ that contains all nodes at distance at most $r$ from $v$ and all edges that have at least one endpoint at distance at most $r-1$ from $v$.
@@ A Generalized Locally Optimal Problem (GLOP) is a pair $(\Pi,\Psi)$ where:
Let $(p_{1},X_{1},Y_{1}),\ldots,(p_{\deg(v)},X_{\deg(v)},Y_{\deg(v)})$ be the labels assigned by $\ell$ to the half-edges incident to $v$. The values $p_{1},\ldots,p_{\deg(v)}$ are called ports of $v$. If $\exists i,j~{}:~{}X_{i}\neq X_{j}\text{ or }Y_{i}\neq Y_{j}$, then $f((G_{r}(v),v))$ is undefined. Let $X=X_{1}$ and $Y=Y_{1}$. Consider a BFS visit on $G_{r}(v)$ where, the order in which the neighbors of some node $u$ are visited, is according to the ports of $u$. If $u$ has ports with values not in $[\deg(u)]$, or two incident ports with the same value, then $f$ is undefined. The labeling $f$ on $(G_{r}(v),v)$ is then obtained by assigning, to the $j$-th visited half-edge, the $j$-th element of $X$, and to the $j$-th visited node, the $j$-th element of $Y$.
@@
[…]
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=M\text{ otherwise,}$,
@@ 3.1 Preliminary definitions and results
A key notion in our algorithm is that of an improving set. As already described in the introduction, the main insight of our approach is that if a node $u$ is part of some sequential flipping process that obtains a globally correct solution, we can detect this in $u$’s $O(\operatorname{log}n)$ neighborhood. What we will find in $u$’s neighborhood is an improving set with a good improving ratio. We will later proof this formally, so to state our algorithm, we give the necessary definitions here.
@@
Let $(\Pi,\Psi)$ be an LOP as defined in Definition 2.7. Consider any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$ not necessarily satisfying the constraints of $\Pi$. Let $\ell_{1}$ denote the labeling of $G$. For a subset $A\subseteq V$ of nodes, we denote by relabeling of $A$ a new output label assignment to the nodes of $A$ and to their incident half-edges. Consider any subset $A$ that can be relabeled, obtaining a new labeling $\ell_{2}$ of $G$, such that $\operatorname{Pot}(G,\ell_{2})<\operatorname{Pot}(G,\ell_{1})$. We say that the pair $(A,\ell_{2})$ is an improving set of $G$ with respect to $\ell_{1}$. The improvement of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the difference $\operatorname{Imp}(A,\ell_{1},\ell_{2}):=\operatorname{Pot}(G,\ell_{1})-\operatorname{Pot}(G,\ell_{2})$. The improving ratio (IR) of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the ratio $\operatorname{IR}(A,\ell_{1},\ell_{2}):=\operatorname{Imp}(A,\ell_{1},\ell_{2})/\absolutevalue{A}$ | if $|A|>0$. For an empty set $A$ of nodes we define $\operatorname{IR}(A,\ell_{1},\ell_{2}):=0$, for any $\ell_{1},\ell_{2}$.
@@
An improving set $(A,\ell_{2})$ is said to be minimal if there is no improving set $(A^{\prime},\ell_{2}^{\prime})$ such that $A^{\prime}\subsetneq A$ and $\operatorname{IR}(A^{\prime},\ell_{1},\ell_{2}^{\prime})>\operatorname{IR}(A,\ell_{1},\ell_{2})$.
Notice that every improving set $(A,\ell_{2})$ contains a minimal improving set, as one can iteratively pick a subset of nodes and a new labeling that increases the improving ratio until no such subset exists. Note that this process must terminate with at least one node, as the empty graph has improving ratio 00.
From a top-down view, our algorithm does nothing else, then relabeling a lot of improving sets. So the changes that our algorithm does to the labeling can be fully described as a sequence of improving sets.
@@
  • For every $1\leq i\leq k$, the pair $(A_{i},\ell_{i})$ is an improving set of $G$ w.r.t. the labeling $\ell_{i-1}$.
@@
In this paragraph, we define and describe the MPX clustering algorithm, which is a key component of our algorithm, following the paper that first introduced it [miller-peng-xu-2013-parallel-graph-decompositions-using].
@@
Lemma 3.5 (Lemma 4.4 of [miller-peng-xu-2013-parallel-graph-decompositions-using]).
@@
Lemma 3.6 ([miller-peng-xu-2013-parallel-graph-decompositions-using]).
@@
Consider any node $v\in V$, and take the node $u$ that realizes $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. By contradiction, suppose there exists a node $w$ such that $\operatorname{dist}_{\text{shift}}(u,u)>\operatorname{dist}_{\text{shift}}(w,u)$. Then, we have that $-\lambda_{u}>\operatorname{dist}(w,u)-\lambda_{w}$. This implies that $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}>\operatorname{dist}(u,v)+\operatorname{dist}(u,w)-\lambda_{w}\geq\operatorname{dist}(w,v)-\lambda_{w}=\operatorname{dist}_{\text{shift}}(w,v)$ by the triangular inequality. Hence, $u$ does not realize the minimum of $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$, reaching a contradiction.
@@
Hence, $\operatorname{dist}_{\text{shift}}(u,v^{\prime})=\operatorname{dist}(u,v^{\prime})+\lambda_{u}\leq\operatorname{dist}(u,v)+\operatorname{dist}(v,v^{\prime})+\lambda_{u}\leq k+\operatorname{dist}_{\text{shift}}(u,v)$. At the same time, for any other $u^{\prime}\neq u$, $\operatorname{dist}_{\text{shift}}(u^{\prime},v^{\prime})=\operatorname{dist}(u^{\prime},v^{\prime})+\lambda_{u^{\prime}}\geq\operatorname{dist}(u^{\prime},v)-\operatorname{dist}(v,v^{\prime})+\lambda_{u^{\prime}}\geq\operatorname{dist}_{\text{shift}}(u^{\prime},v)-k$. Now notice that $\Pr(\cup_{u^{\prime}\neq u}\{\absolutevalue{\operatorname{dist}_{\text{shift}}(u,v)-\operatorname{dist}_{\text{shift}}(u^{\prime},v)}\leq k\})<1/2$ by Lemma 3.5, hence the thesis.

### v3

# Distributed Algorithms for Potential Problems

@@ 1.1 Locally optimal cuts
We say that a cut is locally optimal if for every node at least half of its incident edges are cut edges. Put otherwise, changing the label of $v$ from $\ell(v)$ to $\ell(v)$ does not increase the size of the cut. Such a cut always exists and is trivial to find in the centralized setting: start with any labeling, find any node that violates the constraint (we call these unhappy nodes) and change its label, and repeat until all nodes are happy. Note that changing the label of an unhappy node $v$ will make node $v$ happy but it might make some of its neighbors unhappy. However, this process has to converge after at most $|\absolutevalue{E}$ steps, since the size of the cut increases in each step by at least one.
@@
While it is trivial to find a locally optimal cut with a centralized algorithm, in this work we are interested in distributed algorithms for finding a locally optimal cut. We work in the usual LOCAL model [Lin92, Pel00] of distributed computing: Each node $v\in V$ represents a computer and each edge $\{u,v\}$ represents a communication link between two computers. Initially each node is only aware of its own unique identifier, its degree, and the total number of nodes $n$. Computation proceeds in synchronous rounds; in each round each node can send a message to each neighbor, receive a message from each neighbor, and update its own state. Eventually each node $v$ has to stop and output its own part of the solution, here $\ell(v)$. The running time of the algorithm is $T(n)$ if in any $n$-node graph all nodes stop after at most $T(n)$ rounds.
@@
For any fixed $\Delta$, the locally optimal cut problem is an example of a locally checkable labeling problem (LCL). LCL problems were first defined by [NS95], and there is a large body of work that has developed a complexity theory of LCL problems in the LOCAL model [CV86, Nao91, Lin92, NS95, BFH+16, FG17, GHK18, BHK+18, CP19, CKP19, RG20, BBO+20, BBO+21, DDL+23, Cha20, BBO+21, GRB22, Suo20].
Most natural LCL problems are by now very well-understood, especially for the deterministic LOCAL model. Indeed, for some restricted families of LCLs and restricted families of graphs, there is now a complete classification of the round complexity of all such problems, see e.g. [BBE+20, CSS23, BHK+17, BBC+22, BBC+19]. For example, there is a broad class of symmetry-breaking problems, all of which have round complexity $\Theta(\operatorname{log}^{*}n)$ [CV86, GPS88, Lin92, Nao91], and many other problems [GHK+20] have turned out to be asymptotically as hard as the sinkless orientation problem [BFH+16], which has round complexity $\Theta(\operatorname{log}n)$ in the deterministic LOCAL model [BFH+16, CKP19, GS17].
Locally optimal cut is perhaps the most elementary LCL problem whose distributed complexity is currently wide open. We do not even know if it admits any nontrivial algorithm with round complexity $o(n)$, not even for the smallest nontrivial degree bound $\Delta=3$. (The case of $\Delta=2$ is fully understood, as it is in essence equivalent to the weak-$2$-coloring problem, which can be solved in $O(\operatorname{log}^{*}n)$ rounds, and this is also tight, see e.g. [BHO+19].)
@@
What makes locally optimal cuts particularly intriguing is that many commonly-used lower bound techniques fail. In particular, one cannot use existential graph-theoretic arguments (in the spirit of e.g. the lower bound for coloring trees in [Lin92] or the lower bound for approximate coloring in [CDG+24]), since a locally optimal cut always exists in any graph.
It is, however, possible to construct a nontrivial reduction that connects locally optimal cuts with the sinkless orientation problem [BHL+19], and this gives the following lower bounds: we need at least $\Omega(\operatorname{log}n)$ rounds in the deterministic LOCAL model, and at least $\Omega(\operatorname{log}\operatorname{log}n)$ rounds in the randomized LOCAL model.
This is essentially everything that is known about locally optimal cuts before this work. There are no upper bounds better than the trivial $O(n)$. Furthermore, the lower bounds do not hold in stronger models such as quantum-LOCAL [GKM09, AF14, ACD+25, BBC+25]—there we do not even know if the true complexity of the problem is closer to $O(1)$ or $\Omega(n)$.
We note that locally optimal cuts have been studied e.g. in the context of descriptive combinatorics also under the names of unfriendly colorings and unfriendly partitions [SM12, Con14, CMU20, CT20].
@@ 1.2 Local potential problems
For example, in the locally optimal cut problem, our node labels are $\{-1,+1\}$. A node is “happy” if at least half of its incident edges are cut edges (this is a local definition in the sense that it only depends on the constant-radius neighborhood of the node). We assign local potential $1$ to each cut edge and potential 0 to all other edges (this is a local potential in the sense that it only depends on the constant-radius neighborhood of the edge). Now we can always turn unhappy nodes into happy nodes by flipping their label, and this will improve the total potential around the node.
@@
and the task is to maximize the sum of local utilities. For example, the maximum cut problem is such an optimization problem; the local utility function assigns value $1$ to each cut edge and 0 to all other edges.
@@
Conversely, super-polylogarithmic lower bounds on the distributed round complexity now directly imply that a problem cannot be represented with the help of a local potential. To give a simple example, stable matchings require a linear number of rounds, even in bounded-degree bipartite graphs [FKP+10], so this problem cannot be represented with local potential functions. Results of this form may be of interest also in a much broader context outside distributed computing.
@@
Key examples of local potential problems include defective coloring problems, see e.g. [BE13, Chapter 6], [CPS17, FG17]. Assume that we have $c$ colors and we want to have defect at most $d$, that is, each node can have at most $d$ neighbors with the same color. If $c(d+1)>\Delta$, this is a local potential problem. Here the potential function is the number of properly colored edges. If a node is unhappy (it has more than $d$ neighbors with the same color), it can find a color that makes it happy. Our algorithm can be used to solve this problem in for any constants $c$, $d$ and $\Delta$ with $c(d+1)>\Delta$ in polylogarithmic number of rounds.
@@ 1.3 New techniques and key ideas
Our algorithm uses the MPX clustering algorithm [MPX13, EN22] to divide the nodes into polylogarithmic-radius clusters with small boundaries. Each cluster can make nodes far from boundaries safe; we can simply compute for each cluster the result of repeatedly finding for each internal node an improving set $A$, and flipping $A$ until all internal nodes are safe.
@@
Our algorithm is randomized, but it can be derandomized (with polylogarithmic cost) by using the standard techniques [GHK18, RG20].
@@ 2.1 Notation and definitions
In this paper, we consider simple undirected graphs unless otherwise stated. For a graph $G=(V,E)$, we denote its set of nodes by $V$ and its set of edges by $E$. If $G$ is not clear from the context we use $V(G)$ to denote the set of nodes of $G$ and $E(G)$ to denote the set of edges of $G$. For any two nodes $u,v\in V(G)$ the distance between $u$ and $v$ is the number of hops in a shortest path between $u$ and $v$, and is denoted by $\operatorname{dist}(u,v)$. For any node $v\in V(G)$, the radius-$r$ neighborhood $\mathcal{N}_{r}[v]$ of $v$ is the set of all nodes at distance at most $r$ from $v$. Let $A\subseteq V(G)$ be a subset of nodes, and $B\subseteq E(G)$ a subset of edges. By $\mathcal{N}_{r}[A]$ we denote the set of all nodes at distance at most $r$ from any node in $A$, and by $\mathcal{N}_{r}[B]$ we denote the set of all nodes that are endpoints of edges in $B$ or are at distance at most $r$ from an endpoint of an edge in $B$. For any two graphs $G,H$, the intersection between $G$ and $H$ is the graph $G\cap H=(V(G)\cap V(H),E(G)\cap E(H))$, and the union between $G$ and $H$ is the graph $G\cup H=(V(G)\cup V(H),E(G)\cup E(H))$. The difference between two graphs $G$ and $H$ is the graph $G\setminus H=(V(G),E(G)\setminus E(H))$.
@@
In this paper, we consider the LOCAL model of distributed computation [Lin92]. We are given a distributed network of $n$ nodes/processors, each of which is capable of arbitrary (but finite) computation. Nodes are connected by edges which represent communication links, and are able to send messages of arbitrarily large (but finite) size to each other: note that nodes can discriminate between different communication links through ports, which are numbered from $1$ to $\Delta$, where $\Delta$ is the maximum degree of the network, that are adversarially assigned. Each node is also given a unique identifier in the set $n^{c}]$ for some constant $c\geq 1$, which is, again, adversarially assigned. The communication network is represented by a simple undirected graph $G=(V,E)$ where $V$ is the set of nodes/processors and $E$ is the set of edges/communication links. All nodes run the same algorithm and computation proceeds in synchronous rounds. In each round, each node sends messages to its neighbors, receives messages from all of its neighbors, and performs local computation. The input of a problem is given in the form of a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of labels. The output of a node is an assignment of labels to the node itself and/or to its incident half-edges: note that this information can be stored in the node through the use of port numbers. When a node determines its own output, we say that the node terminates its computation. The running time of a LOCAL algorithm is the number of communication rounds it takes for all nodes to produce their output. We also refer to the running time of an algorithm as the locality of the algorithm. The complexity of a problem is the minimum running time (across all possible algorithms) needed to solve the problem. We can also assume that nodes are equipped with some source of randomness (independently of other nodes), such as an arbitrarily long random bit string, which they can use during the computation. In this case, we talk about the randomized LOCAL model as opposed to the deterministic LOCAL model. In case of randomized computation, we require that the problem is solved with probability at least 11/n1-1/n.
@@ 2.2 Problems of interest
Let $H=(V,E)$ be any graph, and let $v\in V$ be any node of $H$. The pair $(H,v)$ is said to be a centered graph and $v$ is said to be the center of $(H,v)$. We say that the pair $(H,v)$ has radius $r$ if all nodes in $H$ are at distance at most $r$ from $v$ and all edges in $H$ have at least one endpoint at distance at most r1r-1 from $v$.
@@
In the following, we define what it means to satisfy a set of constraints. For a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ and a node $v\in V$, we denote by $G_{r}(v)$ the $(\mathcal{V},\mathcal{E})$-labeled subgraph of $G$ that contains all nodes at distance at most $r$ from $v$ and all edges that have at least one endpoint at distance at most r1r-1 from $v$.
@@ A Generalized Locally Optimal Problem (GLOP) is a pair $(\Pi,\Psi)$ where:
Let $(p_{1},X_{1},Y_{1}),\ldots,(p_{\deg(v)},X_{\deg(v)},Y_{\deg(v)})$ be the labels assigned by $\ell$ to the half-edges incident to $v$. The values $p_{1},\ldots,p_{\deg(v)}$ are called ports of $v$. If $\exists i,j~:~X_{i}\neq X_{j}\text{ or }Y_{i}\neq Y_{j}$, then $f((G_{r}(v),v))$ is undefined. Let $X=X_{1}$ and $Y=Y_{1}$. Consider a BFS visit on $G_{r}(v)$ where, the order in which the neighbors of some node $u$ are visited, is according to the ports of $u$. If $u$ has ports with values not in $\deg(u)]$, or two incident ports with the same value, then $f$ is undefined. The labeling $f$ on $(G_{r}(v),v)$ is then obtained by assigning, to the $j$-th visited half-edge, the $j$-th element of $X$, and to the $j$-th visited node, the $j$-th element of $Y$.
@@
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=\Psi_{g}(G_{r}(v),v)\text{ if $g$ on $(G_{r^{\prime}=2r}(v),v)$ is defined, and}$
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=M\text{ otherwise,}$
@@ 3.1 Preliminary definitions and results
A key notion in our algorithm is that of an improving set. As already described in the introduction, the main insight of our approach is that if a node $u$ is part of some sequential flipping process that obtains a globally correct solution, we can detect this in $u$’s $O(\operatorname{log}n)$ neighborhood. What we will find in $u$’s neighborhood is an improving set with a good improving ratio. We will later prove this formally, so to state our algorithm, we give the necessary definitions here.
@@
Let $(\Pi,\Psi)$ be an LOP as defined in Definition 2.7. Consider any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$ not necessarily satisfying the constraints of $\Pi$. Let $\ell_{1}$ denote the labeling of $G$. For a subset $A\subseteq V$ of nodes, we denote by relabeling of $A$ a new output label assignment to the nodes of $A$ and to their incident half-edges. Consider any subset $A$ that can be relabeled, obtaining a new labeling $\ell_{2}$ of $G$, such that $\operatorname{Pot}(G,\ell_{2})<\operatorname{Pot}(G,\ell_{1})$. We say that the pair $(A,\ell_{2})$ is an improving set of $G$ with respect to $\ell_{1}$. The improvement of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the difference $\operatorname{Imp}(A,\ell_{1},\ell_{2}):=\operatorname{Pot}(G,\ell_{1})-\operatorname{Pot}(G,\ell_{2})$. The improving ratio (IR) of $(A,\ell_{2})$ with respect to $\ell_{1}$ is defined as the ratio $|\operatorname{IR}(A,\ell_{1},\ell_{2}):=\operatorname{Imp}(A,\ell_{1},\ell_{2})/\absolutevalue{A}$ if $|A|>0$. For an empty set $A$ of nodes we define $\operatorname{IR}(A,\ell_{1},\ell_{2}):=0$, for any $\ell_{1},\ell_{2}$. By diameter of an improving set $(A,\ell_{A})$, we denote the weak diameter of the subgraph $G[A]$ induced by $A$.
@@
An improving set $(A,\ell_{2})$ is said to be minimal if there is no improving set $(A^{\prime},\ell_{2}^{\prime})$ such that $A^{\prime}\subseteq A$ and $\operatorname{IR}(A^{\prime},\ell_{1},\ell_{2}^{\prime})>\operatorname{IR}(A,\ell_{1},\ell_{2})$.
Notice that every improving set $(A,\ell_{2})$ contains a minimal improving set, as one can iteratively pick a subset of nodes (possibly, the same set) and a new labeling that increases the improving ratio until no such subset exists. Note that this process must terminate with at least one node, as the empty graph has improving ratio 0.
From a top-down view, our algorithm does nothing else than relabeling a lot of improving sets. So the changes that our algorithm does to the labeling can be fully described as a sequence of improving sets.
@@
  • For every $1\leq i\leq k$, the pair $(A_{i},\ell_{i})$ is a minimal improving set of $G$ w.r.t. the labeling $\ell_{i-1}$.
@@
In this paragraph, we define and describe the MPX clustering algorithm, which is a key component of our algorithm, following the paper that first introduced it [MPX13].
@@
Lemma 3.5 (Lemma 4.4 of [MPX13]).
@@
Lemma 3.6 ([MPX13]).
@@
Consider any node $v\in V$, and take the node $u$ that realizes $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. By contradiction, suppose there exists a node $w$ such that $\operatorname{dist}_{\text{shift}}(u,u)>\operatorname{dist}_{\text{shift}}(w,u)$. Then, we have that $\lambda_{u}>\operatorname{dist}(w,u)-\lambda_{w}$. This implies that $\operatorname{dist}_{\text{shift}}(u,v)=\operatorname{dist}(u,v)-\lambda_{u}>\operatorname{dist}(u,v)+\operatorname{dist}(u,w)-\lambda_{w}\geq\operatorname{dist}(w,v)-\lambda_{w}=\operatorname{dist}_{\text{shift}}(w,v)$ by the triangular inequality. Hence, $u$ does not realize the minimum of $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$, reaching a contradiction.
Now, we say that each node $v$ joins the cluster of the node $u$ that minimizes the shifting distance from $u$ to $v$, that is, $\min_{v^{\prime}\in V}\operatorname{dist}_{\text{shift}}(v^{\prime},v)$. Let $d=O(\operatorname{log}n/\rho)$. In [MPX13], it is shown that this procedure produces a $(\rho,d)$-decomposition of $G$ with high probability. While the original procedure is defined for the PRAM model, [EN22] showed that it can be implemented in the LOCAL model through an $O(\operatorname{log}n/\rho)$-round randomized LOCAL algorithm.